# Fractal Graph Contrastive Learning

### v2

@@ Abstract
While Graph Contrastive Learning (GCL) has attracted considerable attention in the field of graph self‑supervised learning, its performance heavily relies on data augmentations that are expected to generate semantically consistent positive pairs. Existing strategies typically resort to random perturbations or local structure preservation, yet lack explicit control over global structural consistency between augmented views. To address this limitation, we propose Fractal Graph Contrastive Learning (FractalGCL), a theory-driven framework that leverages fractal self-similarity to enforce global topological coherence. FractalGCL introduces two key innovations: a renormalisation-based augmentation that generates structurally aligned positive views via box coverings; and a fractal-dimension-aware contrastive loss that aligns graph embeddings according to their fractal dimensions. While combining the two innovations markedly boosts graph-representation quality, it also adds non-trivial computational overhead. To mitigate the computational overhead of fractal dimension estimation, we derive a one-shot estimator by proving that the dimension discrepancy between original and renormalised graphs converges weakly to a centred Gaussian distribution. This theoretical insight enables a reduction in dimension computation cost by an order of magnitude, cutting overall training time by approximately 61%. The experiments show that FractalGCL not only delivers state-of-the-art results on standard benchmarks but also outperforms traditional baselines on traffic networks by an average margin of about remarkably 7%. Codes are available at (https://anonymous.4open.science/r/FractalGCL-0511/).
@@ 1 Introduction
Graph contrastive learning (GCL) has emerged as a popular self-supervised paradigm for graph representation learning hu2020strategies ; xia2022simgrace ; joao ; gclsurvey1 ; gclsurvey2 ; xu2018powerful . By forcing models to discriminate positive pairs from negative pairs, it alleviates the knowledge scarcity problems gsslsurvey1 ; gsslsurvey2 ; graphscarcity1 ; graphscarcity2 , and it also serves as an effective pretext task for pre-training graph foundation models gfm1 ; gfm2 . As graphs possess non-Euclidean topology, researchers must tailor contrastive learning frameworks to graph-specific properties. Therefore, GCL has formed unique lines of research, which cover stages including augmenting graph data gclda1 ; gclda2 ; gclda3 , designing contrastive modes gclmode1 ; gclmode2 ; gclmode3 , and optimizing contrastive objectives gclopt1 ; gclopt2 ; gclopt3 .
Among current studies, data augmentation remains a pivotal challenge in graph contrastive learning, as the quality of positive and negative sample pairs fundamentally determines the capacity of a graph model to extract meaningful knowledge and the quality of learned representations. While negative samples are typically generated by contrasting views from structurally distinct graphs or subgraphs graphcl , which ensures divergent distributional characteristics, the generation of semantically coherent positive samples remains a critical bottleneck. Specifically, existing approaches often rely on random perturbations (e.g., node/edge deletion, attribute masking) or fixed topological constraints (e.g., hierarchy preservation), which provide only incomplete guarantees for maintaining structural consistency. These methods lack an explicit mechanism to ensure global similarity between the original graph and its augmented views, leading to potential mismatches in semantic alignment. This gap naturally raises a fundamental question: Can we design a principled graph-level criterion to enforce global structural consistency during positive sample generation?
This critical question directs attention to a fundamental yet often overlooked global property of graphs—their inherent self‑similarity and hierarchical complexity, which is mathematically formalised through the concept of fractal. Fractal geometry edgar2008measure ; mandelbrot1983fractal ; mandelbrot1989fractal is a field of mathematics that explores irregular shapes whose intricate detail persists across different scales, appearing in patterns such as snowflakes, coastlines, and branching trees. Fractal graphs are networks that possess fractal properties, effectively transplanting fractal concepts from Euclidean space onto graph structures (see Figure 1). Given the prevalence of fractal graphs in natural and society, their fractal properties likely play a significant yet under-explored role in improving graph representations via GCL.
@@
Figure 1: An example of evolving theoretical fractal graph neroli2024fractal
@@
  • From Theory to Implementation. To further optimize Fractal GCL for practical implementation, we conducted a series of theoretical analyses to prove that the gap between the original and renormalised box dimensions converges weakly to a centred Gaussian measure, enabling a lightweight approach that markedly accelerates training.
@@ 2 Preliminary Experiments
Preliminary experiment 1
We assessed how well each graph in six graph classification benchmarks follows fractal (power-law) scaling by fitting a log–log box-counting linear regression and recording its coefficient of determination $R^{2}$: the closer $R^{2}$ is to 1, the more convincingly the graph is fractal.
@@
Preliminary experiment 2
We augmented a new feature box dimension, a type of fractal dimension introduced in Section 3.1, to original features in each benchmark, and report the difference in classification accuracy before and after augmentation. Remarkably, four out of six benchmarks show statistically significant gains ($p<0.05$), strongly indicating that fractal information captures unique topological features that previous models fail to capture. Besides, the largest improvement ($p=9.7\times 10^{-4}$) appears on the most fractal and large-scale graphs, suggesting that box dimension is especially informative when global self-similarity is pronounced.
@@ 3.1 Background of Fractal Geometry
Let $G$ be an infinite graph equipped with the graph distance $d_{G}$. An $L$ box covering of $G$ is a collection of subgraphs $\{U_{i}\}_{i\in\mathcal{I}}$, indexed by $\mathcal{I}$, such that $\bigcup_{i\in\mathcal{I}}U_{i}=G,$ and $\operatorname{diam}(U_{i})\leq L$ for any $i\in\mathcal{I}.$ Denote by $N_{L}(G)$ the minimum number of subgraphs required for an $L$ box covering of $G$. The Minkowski dimension (or box dimension) of $G$ is then defined as
$\dim_{B}(G):=\lim_{L/\operatorname{diam}(G)\to 0}\frac{\log N_{L}(G)}{-\,\log\bigl{(}L/\operatorname{diam}(G)\bigr{)}},$
provided the limit exists. If $0<\dim_{B}(G)<\infty$, we say that $G$ exhibits a fractal property and call $G$ a fractal (Minkowski) graph.
@@ 3.2 New Augmentation: Graph Renormalisation
In contrastive learning, constructing an “augmented graph” that retains structural similarity while preserving appropriate differences from the original is crucial. Here, we introduce the concept of the renormalisation graph, whose core idea has appeared in various literature (e.g., in multi-scale network analysis and abstractions of complex networks song2005self ), but whose application to augmentations for contrastive learning is relatively novel.
@@
Let $G$ be a given graph, and let $\{U_{i}\}_{i\in\mathcal{I}}$ be an $L$ box-covering of $G$. We construct the renormalised graph $\mathscr{R}(G)$ as follows: (i) Collapse each covering set $U_{i}$ into a single supervertex $v_{i}$. (ii) If there exists at least one edge in $G$ connecting a vertex in $U_{i}$ to a vertex in $U_{j}$ (with $i\neq j$), then place a superedge between $v_{i}$ and $v_{j}$ in $\mathscr{R}(G)$.
The resulting graph $\mathscr{R}(G)$ is called the renormalised graph of $G$ at the given scale $L$.
@@
For any Minkowski (box-dimensional) graph $G$, mathematically,
$\dim_{B}\bigl{(}\mathscr{R}(G)\bigr{)}=\dim_{B}(G).$
@@
Theorem 3.3 formally states that the box dimension is invariant under renormalisation, ensuring that $\mathscr{R}(G)$ and $G$ share the same intrinsic fractal complexity.
@@ 3.3 Novel Loss: Fractal-Dimension Based
Mapping from Graph $G$ to Representation $\mathbf{z}$
we apply $\mathscr{R}$ to obtain the augmented graph $\mathscr{R}(G_{n})$. We then use a GNN-based encoder $f_{\theta}(\cdot)$ and a readout function $\mathrm{Readout}(\cdot)$ to produce a graph-level embedding, and finally apply a projection head $g_{\phi}(\cdot)$ to map it into the contrastive space: $\mathbf{z}_{n}:=g_{\phi}\Bigl{(}\mathrm{Readout}\bigl{(}f_{\theta}(\mathscr{R}(G_{n}))\bigr{)}\Bigr{)}.$
Contrastive Loss with Fractal Weight
We define the fractal dimension discrepancy weight between $G_{n}$ and its augmented version $\mathscr{R}(G_{n})$ as
$\varphi\bigl{(}G_{n},\mathscr{R}(G_{n})\bigr{)}:=\exp\Bigl{(}\alpha\,\bigl{|}\dim_{B}\bigl{(}G_{n}\bigr{)}-\dim_{B}\bigl{(}\mathscr{R}(G_{n})\bigr{)}\bigr{|}\Bigr{)},$
where $\dim_{B}(\cdot)$ denotes the (estimated) fractal dimension of a graph, and $\alpha\geq 0$ is a scaling factor.
@@
Denote the similarity by $s:=\mathrm{sim}\bigl{(}\mathbf{z}_{G},\mathbf{z}_{\mathscr{R}(G)}\bigr{)}$. Keeping all other batch terms fixed, the fractal-weighted InfoNCE loss $\ell_{\mathrm{fractal}}$ satisfies
$\Bigl{\lvert}\partial\ell_{\mathrm{fractal}}/\partial s\Bigr{\rvert}=w\,\Bigl{\lvert}\partial\ell_{\mathrm{InfoNCE}}/\partial s\Bigr{\rvert},w=\exp(\alpha\Delta)\text{ increases strictly with }\Delta.$
@@
Intuitively, a larger fractal‐dimension gap $\Delta$ amplifies the positive-pair gradient, forcing the model to pull the two views closer, while $\Delta=0$ reduces to the ordinary InfoNCE case. The lemma therefore formalises how the weight $\exp(\alpha\Delta)$ adaptively injects fractal similarity into the optimisation dynamics.
@@
If $\Delta_{1}<\Delta_{2}$ and $s_{1}-s_{2}\leq\tau\,\alpha\bigl{(}\Delta_{2}-\Delta_{1}\bigr{)}$, then the fractal-weighted InfoNCE losses satisfy
@@ 3.4 Computational Dilemma and Its Solution
For the greedy box-covering procedure in Algorithm 3, the worst-case running time $T(V)$ obeys
$\Omega\bigl{(}V^{2}\bigr{)}\;\leq\;T(V)\;\leq\;O\bigl{(}V^{3}\bigr{)}.$
@@
A Theoretical Solution to the Dimension Dilemma
We avoid the heavy cost of recomputing $\dim\bigl{(}\mathscr{R}(G)\bigr{)}$ at every augmentation step by modelling the finite‐size deviation $\Delta(G):=\dim(G)-\dim\bigl{(}\mathscr{R}(G)\bigr{)}$ as a random perturbation whose variance vanishes as the graph grows. The argument proceeds in three succinct steps.
Step 1. Finite-diameter error magnitude. Let $G$ be a graph with diameter $D:=\operatorname{diam}(G)$. Write $\hat{m}_{G}$ for the OLS slope used to estimate $\dim(G)$ and $\sigma^{2}$ for the log–residual variance.
@@ A Theoretical Solution to the Dimension Dilemma
For any graph $G$, $\operatorname{SE}(\hat{m}_{G})\;\sim\;\sqrt{6}\,\sigma\bigl{[}\sqrt{D}\,\log D\bigr{]}^{-1}.$
@@
See Appendix A.5. ∎
Lemma 3.7 quantifies how rapidly the slope uncertainty shrinks: the error decays as $1/\bigl{(}\sqrt{D}\,\log D\bigr{)}$.
@@
$\sqrt{D}\,(\hat{m}_{G}-m_{G})\xrightarrow{\mathscr{D}}\mathcal{N}(0,\sigma^{2}),\qquad\sqrt{D}\,(\hat{m}_{\mathscr{R}}-m_{\mathscr{R}})\xrightarrow{\mathscr{}}\mathcal{N}(0,\sigma^{2}).$
@@
Lemma 3.8 states that, once rescaled by $\sqrt{D}$, the slope estimator for either graph becomes asymptotically Gaussian with a diameter-independent variance $\sigma^{2}$. Hence any finite-size fluctuation of the estimated dimension is fully captured by a normal term whose magnitude is controlled only by $D$.
@@
Let $\mu_{G}$ be the probability measure induced by the random variable $\Delta(G)=\dim(G)-\dim\bigl{(}\mathscr{R}(G)\bigr{)}$ on a graph $G$. Under Hypothesis A.6 and with the notation of Lemmas 3.7 and 3.8, we have
[…]
as $D\to\infty$. In particular, $\tau^{2}(D)\to 0$, so the limiting distribution degenerates to the Dirac measure $\delta_{0}$; i.e. $\Delta(G)\to 0$ in probability.
@@
When $D\to\infty$ the variance $\tau^{2}(D)$ vanishes and $\mu_{G}\to\delta_{0}$, so $\dim(G)=\dim\bigl{(}\mathscr{R}(G)\bigr{)}$ with probability $1$ in the infinite-diameter limit.

### v3

@@ Abstract
While Graph Contrastive Learning (GCL) has attracted considerable attention in the field of graph self‑supervised learning, its performance heavily relies on data augmentations that are expected to generate semantically consistent positive pairs. Existing strategies typically resort to random perturbations or local structure preservation, yet lack explicit control over global structural consistency between augmented views. To address this limitation, we propose Fractal Graph Contrastive Learning (FractalGCL), a theory-driven framework introducing two key innovations: a renormalisation-based augmentation that generates structurally aligned positive views via box coverings; and a fractal-dimension-aware contrastive loss that aligns graph embeddings according to their fractal dimensions, equipping the method with a fallback mechanism guaranteeing a performance lower bound even on non-fractal graphs. While combining the two innovations markedly boosts graph-representation quality, it also adds non-trivial computational overhead. To mitigate the computational overhead of fractal dimension estimation, we derive a one-shot estimator by proving that the dimension discrepancy between original and renormalised graphs converges weakly to a centred Gaussian distribution. This theoretical insight enables a reduction in dimension computation cost by an order of magnitude, cutting overall training time by approximately 61%. The experiments show that FractalGCL not only delivers state-of-the-art results on standard benchmarks but also outperforms traditional and latest baselines on traffic networks by an average margin of about remarkably 4%. Codes are available at (https://anonymous.4open.science/r/FractalGCL-0511/).
@@ 1 Introduction
Graph contrastive learning (GCL) has emerged as a popular self-supervised paradigm for graph representation learning (Hu et al., 2020; Xia et al., 2022a; You et al., 2021; Ju et al., 2024; Liu et al., 2022a; Xu et al., 2018). By forcing models to discriminate positive pairs from negative pairs, it alleviates the knowledge scarcity problems (Wu et al., 2021; Xie et al., 2022; Chen et al., 2025; Shi et al., 2025), and it also serves as an effective pretext task for pre-training graph foundation models (Liu et al., 2023; Huang et al., 2024). As graphs possess non-Euclidean topology, researchers must tailor contrastive learning frameworks to graph-specific properties. Therefore, GCL has formed unique lines of research, which cover stages including augmenting graph data (Liu et al., 2022b; Rong et al., 2019; Sun et al., 2021), designing contrastive modes (Ju et al., 2023; Ren et al., 2021; Park et al., 2020), and optimizing contrastive objectives (Hjelm et al., 2018; Xia et al., 2022b; Zhang et al., 2022).
Among current studies, data augmentation remains a pivotal challenge in graph contrastive learning, as the quality of positive and negative sample pairs fundamentally determines the capacity of a graph model to extract meaningful knowledge and the quality of learned representations. While negative samples are typically generated by contrasting views from structurally distinct graphs or subgraphs (You et al., 2020), which ensures divergent distributional characteristics, the generation of semantically coherent positive samples remains a critical bottleneck. Specifically, existing approaches often rely on random perturbations (e.g., node/edge deletion, attribute masking) or fixed topological constraints (e.g., hierarchy preservation), which provide only incomplete guarantees for maintaining structural consistency. These methods lack an explicit mechanism to ensure global similarity between the original graph and its augmented views, leading to potential mismatches in semantic alignment. This gap naturally raises a fundamental question: Can we design a principled graph-level criterion to enforce global structural consistency during positive sample generation?
This critical question directs attention to a fundamental yet often overlooked global property of graphs—their inherent self‑similarity and hierarchical complexity, which is mathematically formalised through the concept of fractal. Fractal geometry (Edgar & Edgar, 2008; Mandelbrot, 1983; 1989) is a field of mathematics that explores irregular shapes whose intricate detail persists across different scales, appearing in patterns such as snowflakes, coastlines, and branching trees. Fractal graphs are networks that possess fractal properties, effectively transplanting fractal concepts from Euclidean space onto graph structures (see Figure 1). Given the prevalence of fractal graphs in natural and society, their fractal properties likely play a significant yet under-explored role in improving graph representations via GCL.
@@
Figure 1: An example of evolving theoretical fractal graph (Neroli, 2024)
@@
  • From Theory to Implementation. To further optimize Fractal GCL for practical implementation, we conducted a series of theoretical analyses to prove that the gap between the original and renormalised box dimensions converges weakly to a centred Gaussian measure, enabling a lightweight approach that markedly accelerates training. We design a principled fallback mechanism, ensuring that even on graphs without fractality our performance is no worse than classical GCL.
@@ 2 Preliminary Experiments
Preliminary experiment 1. We assessed how well each graph in six graph classification benchmarks follows fractal (power-law) scaling by fitting a log–log box-counting linear regression and recording its coefficient of determination $R^{2}$: the closer $R^{2}$ is to 1, the more convincingly the graph is fractal.
@@
Preliminary experiment 2. We augmented a new feature box dimension, a type of fractal dimension introduced in Section 3.1, to the original features in each benchmark, and report the difference in classification accuracy before and after augmentation. Remarkably, four out of six benchmarks show statistically significant gains ($p<0.05$), strongly indicating that fractal information captures unique topological features that previous models fail to capture. Besides, the largest improvement ($p=9.7\times 10^{-4}$) appears on the most fractal and large-scale graphs, suggesting that box dimension is especially informative when global self-similarity is pronounced.
@@ 3.1 Background of Fractal Geometry
Let $G$ be an infinite graph equipped with the graph distance $d_{G}$. An $L$ box covering of $G$ is a collection of subgraphs $\{U_{i}\}_{i\in\mathcal{I}}$, indexed by $\mathcal{I}$, such that $\bigcup_{i\in\mathcal{I}}U_{i}=G,$ and the diameter $\mathrm{diam}(U_{i})\leq L$ for any $i\in\mathcal{I}.$ Denote by $N_{L}(G)$ the minimum number of subgraphs required for an $L$ box covering of $G$. The Minkowski dimension (or box dimension) of $G$ is then defined as
$\dim_{\mathrm{B}}(G):=\lim_{L/\operatorname{diam}(G)\to 0}\frac{\log N_{L}(G)}{-\log\bigl(L/\operatorname{diam}(G)\bigr)},$
provided the limit exists. If $0<\dim_{\mathrm{B}}(G)<\infty$, we say that $G$ exhibits a fractal property and call $G$ a fractal (Minkowski) graph. In network science, it can also be stated as $N_{L}(G)\sim\left(\frac{\operatorname{diam}(G)}{L}\right)^{\beta}$ for $L\ll\operatorname{diam}(G)$.
@@ 3.2 New Augmentation: Graph Renormalisation
In contrastive learning, constructing an “augmented graph” that retains structural similarity while preserving appropriate differences from the original is crucial. Here, we introduce the concept of the renormalisation graph, whose core idea has appeared in various literature (e.g., in multi-scale network analysis and abstractions of complex networks (Song et al., 2005)), but whose application to augmentations for contrastive learning is relatively novel.
@@
Let $G$ be a given graph, and let $\{U_{i}\}_{i\in\mathcal{I}}$ be an $L$ box-covering of $G$. We construct the renormalised graph $\mathscr{R}(G)$ as follows: (i) Collapse each covering set $U_{i}$ into a single supervertex $v_{i}$. (ii) If there exists at least one edge in $G$ connecting a vertex in $U_{i}$ to a vertex in $U_{j}$ (with $i\neq j$), then place a superedge between $v_{i}$ and $v_{j}$ in $\mathscr{R}(G)$. The resulting graph $\mathscr{R}(G)$ is called the renormalised graph of $G$ at the given scale $L$. $\mathscr{R}(G)$ is equipped with the unweighted shortest-path metric induced by its adjacency matrix.
@@
For any Minkowski (box-dimensional) infinite graph $G$, mathematically,
$\dim_{\mathrm{B}}\bigl(\mathscr{R}(G)\bigr)=\dim_{\mathrm{B}}(G).$
@@
Note that for any graph $G$ of infinite diameter, its renormalisation $\mathscr{R}(G)$ necessarily retains infinite diameter; hence the Minkowski dimension remains well-defined. Theorem 3.3 formally states that the box dimension is invariant under renormalisation, ensuring that $\mathscr{R}(G)$ and $G$ share the same intrinsic fractal complexity.
@@ 3.3 Novel Loss: Fractal-Dimension Based
Mapping from Graph $G$ to Representation $\mathbf{z}$. we apply $\mathscr{R}$ to obtain the augmented graph $\mathscr{R}(G_{n})$, where $G_{n}$ is the n-th original graph in a mini-batch. We then use a GNN-based encoder $f_{\theta}(\cdot)$ and a readout function $\mathrm{Readout}(\cdot)$ to produce a graph-level embedding, and finally apply a projection head $g_{\phi}(\cdot)$ to map it into the contrastive space: $\mathbf{z}_{n}:=g_{\phi}\Bigl(\mathrm{Readout}\bigl(f_{\theta}(\mathscr{R}(G_{n}))\bigr)\Bigr).$
Contrastive Loss with Fractal Weight. We define the fractal dimension discrepancy weight between $G_{n}$ and its augmented version $\mathscr{R}(G_{n})$ as
$\varphi\bigl(G_{n},\mathscr{R}(G_{n})\bigr):=\exp\Bigl(\alpha\bigl|\dim_{\mathrm{B}}\bigl(G_{n}\bigr)-\dim_{\mathrm{B}}\bigl(\mathscr{R}(G_{n})\bigr)\bigr|\Bigr),$
where $\dim_{\mathrm{B}}(\cdot)$ denotes the (estimated) Minkowski dimension of a graph, and $\alpha\geq 0$ is a scaling factor.
@@
Denote the similarity by $s(G):=\mathrm{sim}\bigl(\mathbf{z}_{G},\mathbf{z}_{\mathscr{R}(G)}\bigr)$ and $\Delta(G):=\dim_{\mathrm{B}}(G)-\dim_{\mathrm{B}}\mathscr{R}(G)$. Keeping all other batch terms fixed, the fractal-weighted InfoNCE loss $\ell_{\mathrm{fractal}}$ satisfies
$\Bigl\lvert\partial\ell_{\mathrm{fractal}}/\partial s(G)\Bigr\rvert=w\Bigl\lvert\partial\ell_{\mathrm{InfoNCE}}/\partial s(G)\Bigr\rvert,w=\exp(\alpha\Delta(G))\text{ increases strictly with }\Delta(G).$
@@
Intuitively, a larger fractal‐dimension gap $\Delta(G)$ amplifies the positive-pair gradient, forcing the model to pull the two views closer, while $\Delta(G)=0$ reduces to the ordinary InfoNCE case. The lemma therefore formalises how the weight $\exp\big(\alpha\Delta(G)\big)$ adaptively injects fractal similarity into the optimisation dynamics.
@@
If $\Delta(H_{1})<\Delta(H_{2})$ and $s(H_{1})-s(H_{2})\leq\tau\alpha\bigl(\Delta(H_{2})-\Delta(H_{1})\bigr)$, then the fractal-weighted InfoNCE losses satisfy
@@ 3.4 Computational Dilemma and Its Solution
For the greedy box-covering procedure in Algorithm 1, the worst-case running time $T(V)$ obeys $\Omega\bigl(V^{2}\bigr)\;\leq\;T(V)\;\leq\;O\bigl(V^{3}\bigr).$
@@
A Theoretical Solution to the Dimension Dilemma. We avoid the heavy cost of recomputing $\dim_{\mathrm{B}}\bigl(\mathscr{R}(G)\bigr)$ at every augmentation step by modelling the finite‐size deviation $\Delta(G):=\dim_{\mathrm{B}}(G)-\dim_{\mathrm{B}}\bigl(\mathscr{R}(G)\bigr)$ as a random perturbation whose variance vanishes as the graph grows. The argument proceeds in three succinct steps.
Step 1. Finite-diameter error magnitude. Denote the diameter of a graph $G$ by $\mathrm{diam}(G)$. Write $\hat{m}_{G}$ for the OLS slope used to estimate $\dim_{\mathrm{B}}(G)$ and $\sigma^{2}$ for the log–residual variance.
@@
$\operatorname{SE}(\hat{m}_{G})\;\sim\;2\sqrt{6}\sigma\bigl[\sqrt{\mathrm{diam}(G)}\log\mathrm{diam}(G)\bigr]^{-1}.$
@@
See Appendix A.5 for details. ∎
Lemma 3.7 quantifies how rapidly the slope uncertainty shrinks: the error decays as $1/\bigl(\sqrt{\mathrm{diam}(G)}\log\mathrm{diam}(G)\bigr)$.
@@
$\sqrt{\mathrm{diam}(G)}(\hat{m}_{G}-m_{G})\xrightarrow{\mathscr{D}}\mathcal{N}(0,\sigma^{2}),\qquad\sqrt{\mathrm{diam}(G)}(\hat{m}_{\mathscr{R}}-m_{\mathscr{R}})\xrightarrow{\mathscr{}}\mathcal{N}(0,\sigma^{2}).$
@@
Lemma 3.8 states that, once rescaled by $\sqrt{\mathrm{diam}(G)}$, the slope estimator for either graph becomes asymptotically Gaussian with a diameter-independent variance $\sigma^{2}$. Hence any finite-size fluctuation of the estimated dimension is fully captured by a normal term whose magnitude is controlled only by $\mathrm{diam}(D)$.
@@
Let $\mu_{G}$ be the probability measure induced by the random variable $\Delta(G)$ on a graph $G$. Under Hypothesis A.6 and with the notation of Lemmas 3.7 and 3.8, we have
$\mu_{G}\;\xrightarrow{\text{w}}\mathcal{N}\bigl(0,\kappa^{2}(\mathrm{diam}(G))\bigr),\qquad\kappa^{2}(\mathrm{diam}(G)):=6\sigma^{2}[\mathrm{diam}(G)(\log\mathrm{diam}(G))^{2}]^{-1},$
as $\mathrm{diam}(G)\to\infty$. In particular, $\kappa^{2}(\mathrm{diam}(G))\to 0$, so the limiting distribution degenerates to the Dirac measure $\delta_{0}$; i.e. $\Delta(G)\to 0$ in probability.
@@
When $\mathrm{diam}(G)\to\infty$ the variance $\kappa^{2}(\mathrm{diam}(G))$ vanishes and $\mu_{G}\to\delta_{0}$, so $\dim_{\mathrm{B}}(G)=\dim_{\mathrm{B}}\bigl(\mathscr{R}(G)\bigr)$ with probability $1$ in the infinite-diameter limit.